Inhaltsverzeichnis

Lord George Gordon Byron , Hon går i skönhet; Lord George Gordon Byron , Sie geht in Schönheit; Lord George Gordon Byron , She Walks in Beauty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Lord George Gordon Byron , Hon går i skönhet; Lord George Gordon Byron , Sie geht in Schönheit; Lord George Gordon Byron , She Walks in Beauty

2/3

Lord George Gordon Byron ,  Lord George Gordon Byron ,  Lord George Gordon Byron ,
Hon går i skönhet  Sie geht in Schönheit  She Walks in Beauty
tradukita de Theodor Lind  tradukita de Adolf Böttger

Hon går i skönhet, lik en natt, Med himmel klar och stjerners glans; Af ljus och skugga mötas gladt I min och blick, hvad skärast fanns, Så vekt, så skönt, så mildradt, att Dess glans af dagrar blott ej vanns.

Sie geht in Schönheit, gleich der Nacht In wolkenlosem Sternenlicht; Des Schattens und des Lichtes Pracht Eint sich in ihrem Angesicht; Aus dem ein milder Schimmer lacht, Der stets dem grellen Tag gebricht.

She walks in Beauty, like the Night Of cloudless climes and starry skies; And all that’s best of dark and bright Meet in her aspect and her eyes; Thus mellowed to the tender light Which Heaven to gaudy day denies –

En stråle bort, en skugga mer, Och borta vore det behag, Som i den mörka locken ler, Som blickar ur de ljufva drag, Der hvarje tanke klart man ser, Hur ren den går i öppen dag.

Ein Strahl hinweg, ein Schatten mehr, Und fort würd’ auch die Anmut sein, Die aus dem Rabenlockenmeer Die Stirn umglänzt mit sanftem Schein, Wo die Gedanken süß und hehr Verkünden, dass ihr Wohnsitz rein.

One shade the more – one ray the less Had half impaired the nameless grace Which waves in every raven tress Or softly lightens o’er her face Where thoughts serenely sweet express How pure – how dear their dwelling place!

Och panna, ögon, kind och mund Och rosor, löjen - allting der Så ljuft-vältaligt hvarje stund Om hjertats renhet vittne bär, Om frid med allt på jordensrund, Om kärlek, som blott oskuld är.

Und auf der Stirn, dem Wangenpaar, Spricht von dem reinsten Jugendmut So sanft beredt, so ruhig klar Des Lächelns Reiz, der Farben Glut, Von einem Herzen wunderbar, Wo Liebe voller Unschuld ruht.

And on that cheek, and o’er that brow, So soft – so calm – yet eloquent, The smiles that win, the tints that glow, But tell of days in goodness spent – A mind at peace with all below – A Heart – whose love is innocent!

…  …  …

Lord George Gordon Byron , Hon går i skönhet; Lord George Gordon Byron , Sie geht in Schönheit; Lord George Gordon Byron , She Walks in Beauty

3/3

Traduko de la Angla poemo “She Walks in Beauty” de Lord George Gordon Byron (∗1788-01-22 – †1824-04-18) en la Svedan de Theodor Lind.

Arg-1109-2254 (2015-01-17 17:43:44)

Tiu ĉi poem-traduko troviĝas en la retejo http: // runeberg. org/ hebrmel/ sub http: // runeberg. org/ hebrmel/ 0011. html ..

Traduko de la Angla poemo “She Walks in Beauty” de Lord George Gordon Byron (∗1788-01-22 – †1824-04-18) en la Germanan de Adolf Böttger (∗1815-05-21 – †1870-11-16).

Arg-1109-2239 (2015-01-17 17:36:39)

Tiun ĉi poem-germanigon mi, Manfred Retzlaff, trovis en la malnova libro “Byron’s sämmtliche Werke”, presita en la germana urbo Lepsiko (Leipzig) en la jaro 1841. Pri la tradukinto vidu la retejon http: // de. wikipedia. org/ wiki/ Adolf_BГ�ttger .

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1109-2238 (2015-01-17 17:34:42)

Tiu ĉi poemo troviĝas en http: // petercochran. files. wordpress. com/ 2009/ 03/ hebrew_melodies. pdf .

http://www.poezio.net/version?poem-id=1109&version-id=2254,2239,2238